OpenStudy (anonymous):

Put the numbers in order, then find the median: -1/2, 0.81, 8.9, 1.5, 1.75, 1.32 A. 1.5 B. 1.41 C. 8.9

OpenStudy (anonymous):

-1/2 .81 1.32 1.5 1.75 8.9 median is half way between 1.32 and 1.5 (1.32 + 1.5)/2 = (2.82)/2 = 1.41 which is the answer

OpenStudy (anonymous):

thanks :)